# Physics-Guided Multi-Fidelity DeepONet for Data-Efficient Flow Field Prediction

### v1

@@ 2.1.2 Enhanced DeepONet with Positional Encoding
Positional encoding (PE) has been widely used in various deep learning architectures, particularly in transformers and attention-based models, to incorporate spatial or temporal information into the input representations [29, 21]. In the context of DeepONet, the trunk net processes coordinate information which can suffer from spectral bias, limiting its ability to capture high-frequency components and intricate spatio-temporal dependencies [31, 28]. By mapping coordinates to a higher-dimensional space beyond the original $(x,y,t)$ representation, PE is expected to enhance the network’s ability to learn complex flow dynamics and temporal dependencies critical for accurate spatio-temporal flow field prediction.
@@
2.2 Multi-Fidelity DeepONet Framework with Physics-Guided Sampling Strategy
@@ 2.2.1 Multi-Fidelity DeepONet Framework
We propose a multi-fidelity DeepONet (MF-DeepONet) framework that efficiently leverages both low-fidelity and high-fidelity data through knowledge transfer. Our two-phase approach first trains a LF DeepONet on abundant low-fidelity data, then strategically transfers these pre-trained networks (branch, trunk, and merge nets) to the high-fidelity model (HF DeepONet). By freezing the branch and trunk parameters (see Figure 3), we preserve the fundamental flow physics captured during low-fidelity training. During the high-fidelity training phase, only the merge network remains trainable, allowing it to learn the specific refinements necessary for accurate high-fidelity predictions. Importantly, the merge network parameters transferred from Phase 1 serve as initial values that are further optimized using the high-fidelity dataset.
@@
To comprehensively understand the impact of different knowledge transfer strategies, we will explore three distinct transfer learning approaches within our framework. While our primary proposed MF-DeepONet strategy employs fine-tuning (where only the merge network is trainable while branch and trunk networks are frozen), we additionally investigate [14]: (1) full-tuning, where all network components (branch, trunk, and merge) are trainable, allowing complete adaptation to high-fidelity data; and (2) linear probing, the most restrictive approach that freezes all layers except the final layer of the merge network. By comparing these approaches, we demonstrate the strength of our fine-tuning strategy while providing valuable insights into the trade-offs between preserving low-fidelity knowledge and adapting to high-fidelity features for multi-fidelity modeling of flow fields using DeepONet.
2.2.2 Time-Derivative Sampling Strategy
@@
After training the low-fidelity (LF) DeepONet model, we aim to efficiently utilize the available high-fidelity (HF) dataset for training the HF DeepONet. While DeepONet’s point-based architecture already offers flexibility in training with either complete or subsampled spatial points (note that effects of sub-sampling will be investigated in Section 4.2), we propose further enhancing this capability through physics-guided sampling. Specifically, we introduce a time-derivative sampling technique that strategically identifies dynamically important regions in the flow field. This approach leverages the temporal derivative information from the pre-trained LF model to guide the selection of HF training points, potentially enabling more efficient training of HF DeepONet by focusing computational resources on regions with significant flow dynamics.
@@
The complete procedure for selecting HF training points using time-derivative guided sampling is formally described in Algorithm 1. This physics-guided approach first computes the temporal derivatives at each spatial point from the HF dataset using the pre-trained LF model through automatic differentiation (lines 4-6). For each HF spatial location, we calculate a score based on the average magnitude of temporal derivatives predicted by the LF model (lines 7-9). Points with larger temporal derivatives—indicating regions of significant flow dynamics—receive higher sampling probabilities after score transformation and normalization (lines 10-11). Rather than deterministically selecting the highest-scoring points, we employ probabilistic sampling using probability defined by these score values, which introduces beneficial stochasticity while still favoring dynamically important regions. A fraction of training points (determined by ratio $r$) is selected based on this probabilistic approach (line 13), while the remaining points are sampled uniformly to maintain spatial coverage (line 14). By strategically focusing computational resources on dynamically important regions while preserving randomness, this approach can achieve superior prediction accuracy with fewer HF samples compared to random sampling without considering any physics. Figure 4 visually demonstrates our time-derivative guided sampling strategy, showing how high-fidelity training points are strategically selected based on a combination of derivative-weighted probability distributions and uniform random sampling to maintain spatial coverage.
Algorithm 1 Time-derivative guided sampling strategy
@@
12:Among $N$ spatial points in the HF dataset for training HF DeepONet:
13: Sample $N\cdot r$ points using multinomial distribution with probabilities $p(\mathbf{x}_{HF})$
14: Sample remaining points uniformly at random
Figure 4: Time-derivative guided sampling demonstration: background shows probability distribution based on temporal derivatives, with black dots representing points selected using this distribution and white circles showing randomly sampled points for spatial coverage.
@@ 3 Data Generation: Flow Fields with Varying Initial Conditions
The training datasets comprise both LF and HF flow field simulations of incompressible Navier-Stokes equations in a two-dimensional periodic domain $[0,1m]\times[0,1m]$. The numerical solution is obtained using a pseudo-spectral solver that employs fast Fourier transforms for spatial discretization, with a 2/3 dealiasing rule to prevent aliasing errors. The solver implements a semi-implicit time integration scheme where advection terms are treated explicitly while diffusion terms are handled implicitly.
@@
where $var_{1}$ and $var_{2}$ are randomly sampled from uniform distributions in the range $\mathcal{U}[1,3]$. The Reynolds number is fixed at $Re=1,000$ for all simulations. The temporal evolution of the flow field is computed with a time step of $\Delta t=2.5\times 10^{-5}[s]$, solving until $t=1s$. While the solver operates at this fine temporal resolution, the flow field data is saved at a coarser interval of $\Delta t_{train}=5\times 10^{-3}[s]$ for training purposes. This output frequency is selected to capture the relevant flow dynamics while maintaining computational efficiency. Importantly, these simulations feature no external forcing after $t=0$, resulting in gradually dissipating flow fields where vorticity magnitude decreases over time. This dissipative behavior creates challenging temporal dynamics that are particularly difficult to predict accurately as the flow evolves further from its initial state. Throughout this study, we focus on predicting the vorticity field, a derived quantity that captures the local rotation in the fluid and presents a more challenging prediction target than primary variables like velocity components. Figure 5 displays the vorticity field at $t=0.5s$ for four distinct initial conditions. The marked diversity in the flow patterns, resulting from variations in $var_{1}$ and $var_{2}$, underscores the complex nature of the dataset.
@@
Figure 5: Vorticity fields at $t=0.5s$ for four distinct initial conditions, each defined by a different combination of $var_{1}$ and $var_{2}$. This diversity in the observed patterns highlights the complexity and challenges of predicting the flow evolution, particularly as dissipative effects intensify over time.
For the LF datasets, we generate 300 training samples for each of three different grid resolutions: 16×16, 32×32, and 64×64. As illustrated in Figure 6, the grid resolution significantly impacts the accuracy of flow field predictions. The coarsest resolution (16×16) shows severe limitations in capturing flow evolution: while it approximates the macroscopic flow patterns at $t=0.25s$, the flow field becomes notably distorted from $t=0.5s$ onward compared to higher-resolution simulations. This suggests that data from such coarse resolutions might be inadequate or even detrimental in a multi-fidelity framework. However, from 32×32 resolution upward, the simulations begin to capture the essential macroscopic flow properties visible in the HF (128×128) case. This observation suggests that LF data from these intermediate resolutions (32×32 and 64×64) could effectively support HF predictions in our multi-fidelity framework.
Figure 6: Effects of grid resolution on vorticity field evolution with fixed parameters $var_{1}=3$ and $var_{2}=1$ (contours represent vorticity values).
@@ 4 Improvement of DeepONet Architecture for Spatio-Temporal Flow Field Prediction
In this section, we focus on enhancing single-fidelity DeepONet performance using only high-fidelity data (128×128 resolution). We systematically investigate the effects of different merge processes, spatial point sampling approaches, and positional encoding strategies to establish the most effective base architecture. This optimized DeepONet architecture will serve as the foundation for both LF and HF models in our multi-fidelity framework presented in Section 5.
@@ 4.1 Effects of Different Merge Process of Branch/Trunk Nets
We first investigate the effectiveness of different strategies for combining outputs from the branch and trunk networks elaborated in Section 2.1.1. To ensure a fair comparison, we carefully adjust the network architectures to maintain similar numbers of learnable parameters across different configurations (approximately 300K parameters). For the basic dot-product variants, Merge Type 0 with and without bias, we employ more complex branch and trunk networks to match the parameter count of models with Type 1 and 2 networks. Across all DeepONet configurations, the branch network consistently takes two input parameters ($var_{1}$ and $var_{2}$) defining initial conditions, while the trunk network processes three spatial-temporal coordinates ($x$, $y$, and $t$). For detailed network architectures, please refer to Table 1. All models are trained for 800 epochs using the Adam optimizer with an initial learning rate of $10^{-3}$, GeLU activation functions, and minibatch size of 1 on an NVIDIA A6000 GPU. To account for training stochasticity and ensure reproducible results, all experiments in this paper are repeated at least three times with different random seeds, and we report the averaged values of these runs.
@@
The results in Table 1 demonstrate two key findings. First, incorporating merge networks significantly improves prediction accuracy compared to two Type 0 approaches suggested by Lu et al. [16]. While the original dot-product architecture was successful for simpler operator learning tasks in other studies, our results show its limitations in complex flow field predictions: the test error with dot-product operations (both with and without bias) remains above 34, indicating substantial prediction inaccuracies. In contrast, the introduction of merge networks dramatically reduces the test error to 17-23, despite requiring longer training times (approximately 8.5 hours compared to 6.2 hours) due to the additional processing of branch and trunk outputs through the distinct merge network. This stark improvement in accuracy—reducing test error from 34.447 to 17.079, a 50.4% improvement—highlights that additional non-linear processing of combined features is crucial for capturing complex spatio-temporal flow dynamics. The merge network effectively acts as a learnable non-linear mapping that can adapt to intricate relationships between the latent representations from branch and trunk networks, a capability that simple dot-product operations alone cannot provide.
Second, among the enhanced architectures, Type 1 (element-wise multiplication) outperforms Type 2 (concatenation), achieving a test error of 17.079 compared to 23.060. This advantage stems from the structured interaction in Type 1, which more naturally aligns with well-established techniques such as modal decomposition (e.g., proper orthogonal decomposition) and spectral expansions, both of which have demonstrated strong performance in CFD.
@@
    Spectral Methods Perspective: Type 1 approach also aligns with spectral methods, where solutions are represented as weighted sums of basis functions [15]. Here, the trunk network acts as a global basis generator, while the branch network provides coefficients—akin to spectral expansions in Fourier or Laplace methods. Since spectral methods have already demonstrated strong performance in solving the Navier-Stokes equations, enforcing this structured decomposition allows Type 1 to better capture the underlying physics of flow evolution, outperforming the less structured feature fusion in Type 2.
@@
4.2 Effects of Spatial Point Sampling and Application of Automatic Mixed Precision
Building upon the optimal merge network configuration (Type 1) identified in Section 4.1, we investigate the impact of spatial point sampling and automatic mixed precision (AMP) training on model performance. Spatial point sampling refers to our approach of selecting only a subset of available spatial locations from the full 128×128 grid for training, leveraging DeepONet’s inherent capability to operate on arbitrary spatial points without requiring fixed grid structures. We explore three sub-sampling strategies: 16×16, 32×32, and 64×64 randomly selected points from the original 128×128 grid. AMP training refers to the strategic use of different numerical precisions (e.g., float32, float16) during model training to reduce memory usage and computational overhead without compromising model accuracy—a particularly relevant approach for DeepONet due to its intensive computational requirements in processing high-dimensional flow field data (for the details of AMP, please refer to Appendix A). For each sampling configuration, we conduct experiments both with and without AMP training to systematically evaluate their combined effects on computational efficiency and prediction accuracy. All other training settings remain consistent with Section 4.1.
Table 2: Comparison of different spatial sampling strategies with and without AMP training: Type 1 merge network is applied.

### v2

@@ 2.1.2 Enhanced DeepONet with Positional Encoding
Positional encoding (PE) has been widely used in various deep learning architectures, particularly in transformers and attention-based models, to incorporate spatial information into the input representations [45, 33]. In the context of DeepONet, the trunk net processes coordinate information which can suffer from spectral bias, limiting its ability to capture high-frequency components and intricate spatio-temporal dependencies [47, 43]. By mapping coordinates to a higher-dimensional space beyond the original $(x,y,t)$ representation, PE is expected to enhance the network’s ability to learn complex flow dynamics and temporal dependencies critical for accurate spatio-temporal flow field prediction.
@@
  • PE($\mathbf{x,y}$): PE is applied only to the spatial coordinates in the trunk network input, preserving the original temporal coordinate.
@@
2.2 Multi-Fidelity DeepONet Framework with Physics-Guided Subsampling Strategy
@@ 2.2.1 Multi-Fidelity DeepONet Framework
We propose a multi-fidelity DeepONet (MF-DeepONet) framework that efficiently leverages both low-fidelity and high-fidelity data through knowledge transfer. Our two-phase approach first trains an LF DeepONet on relatively abundant low-fidelity data, then strategically transfers these pre-trained networks (LF DeepONet) to the high-fidelity model (HF DeepONet). By freezing the branch and trunk parameters of HF DeepONet (see Figure 3), we preserve the fundamental flow physics captured during low-fidelity training. During the high-fidelity training phase, only the merge network remains trainable, allowing it to learn the specific refinements necessary for accurate high-fidelity predictions. Importantly, the merge network parameters transferred from Phase 1 serve as initial values that are further optimized using the high-fidelity dataset.
@@
To comprehensively understand the impact of different knowledge transfer strategies, we will explore three distinct transfer learning approaches within our framework. While our primary proposed MF-DeepONet strategy employs fine-tuning (where only the merge network is trainable while branch and trunk networks are frozen), we additionally investigate [23]: (1) full-tuning, where all network components (branch, trunk, and merge) of HF DeepONet are trainable, allowing complete adaptation to high-fidelity data; and (2) linear probing, the most restrictive approach that freezes all layers except the final layer of the merge network.
2.2.2 Physics-Guided Subsampling Strategy
@@
After training the LF DeepONet model, we aim to efficiently utilize the available HF dataset for training the HF DeepONet. While DeepONet’s point-based architecture already offers flexibility in training with either complete or subsampled spatial points (note that effects of subsampling will be investigated in Section 4.2), we propose further enhancing this capability through physics-guided subsampling. Specifically, we introduce a time-derivative subsampling technique that strategically identifies dynamically important regions in the flow field. This approach leverages the temporal derivative information from the pre-trained LF model to guide the selection of spatial query points from the HF dataset, potentially enabling more efficient training of HF DeepONet by focusing computational resources on regions with significant flow dynamics.
@@
The complete procedure for selecting HF training points using time-derivative guided subsampling is formally described in Algorithm 1. This physics-guided approach first computes the temporal derivatives at each spatial point from the HF dataset using the pre-trained LF model through automatic differentiation (lines 4-6). For each HF spatial location, we calculate a score based on the average magnitude of temporal derivatives predicted by the LF model (lines 7-9). Points with larger temporal derivatives—indicating regions of significant flow dynamics—receive higher sampling probabilities after score transformation and normalization (lines 10-11). Rather than deterministically selecting the highest-scoring points, we employ probabilistic sampling using probability defined by these score values, which introduces beneficial stochasticity while still favoring dynamically important regions. The final training set of $N_{subsampled}$ points is constructed by first selecting a fraction ($N_{subsampled}\cdot r$) via the probabilistic approach (line 13), and then sampling the remaining points ($N_{subsampled}\cdot(1-r)$) uniformly to ensure spatial coverage (line 14). By strategically focusing computational resources on dynamically important regions while preserving randomness, this approach can achieve superior prediction accuracy by using a more informative, smaller subset of HF spatial training points than a purely random sampling would. Figure 4 visually demonstrates our time-derivative guided subsampling strategy, showing how high-fidelity training points are strategically selected based on a combination of derivative-weighted probability distributions and uniform random sampling to maintain spatial coverage.
Algorithm 1 Time-derivative guided subsampling strategy
@@
12:To subsample $N_{subsampled}$ spatial points from the HF dataset for training HF DeepONet:
13: Sample $N_{subsampled}\cdot r$ points using multinomial distribution with probabilities $p(\mathbf{x}_{HF})$
14: Sample remaining $N_{subsampled}\cdot(1-r)$ points uniformly at random
Figure 4: Time-derivative guided subsampling demonstration: background shows probability distribution based on temporal derivatives, with black dots representing points selected using this distribution and white circles showing randomly sampled points for spatial coverage.
2.2.3 Key Considerations for Effective Multi-Fidelity Modeling
Our MF-DeepONet framework with physics-guided subsampling fundamentally relies on the predictive capability of the pre-trained LF DeepONet, meaning its effectiveness is inherently tied to the quality and computational efficiency of the low-fidelity dataset. Therefore, a critical preliminary step in applying any MF strategy, including ours, is the careful selection and validation of the LF model. This selection process must consider two key aspects: (1) the degree to which the LF model can capture the essential physical phenomena relevant to the HF model, as this dictates the potential for effective knowledge transfer, and (2) the computational cost of generating LF data, since an MF approach achieves its overall efficiency gain primarily when LF simulations are meaningfully less computationally expensive than HF simulations. If the LF data are not significantly less expensive to generate, or if they fail to adequately represent the critical physics of the HF data, the advantages of the multi-fidelity approach are naturally diminished. However, this limitation is not unique to our framework—it represents a common challenge shared across multi-fidelity modeling strategies that incorporate LF data, whether explicitly or implicitly.
@@ 3 Data Generation: Flow Fields with Varying Initial Conditions
The training datasets comprise both LF and HF flow simulations of incompressible Navier-Stokes equations in a two-dimensional periodic domain $[0,1m]\times[0,1m]$. The numerical solution is obtained using a pseudo-spectral solver that employs fast Fourier transforms for spatial discretization, with a 2/3 dealiasing rule to prevent aliasing errors. The solver implements a semi-implicit time integration scheme where advection terms are treated explicitly while diffusion terms are handled implicitly.
@@
where $var_{1}$ and $var_{2}$ are randomly sampled from uniform distributions in the range $\mathcal{U}[1,3]$. The Reynolds number is fixed at $Re=1,000$ for all simulations. The temporal evolution of the flow field is computed with a time step of $\Delta t=2.5\times 10^{-5}[s]$, solving until $t=1s$. While the solver operates at this fine temporal resolution, the flow field data is extracted at a coarser interval of $\Delta t_{train}=5\times 10^{-3}[s]$ for training purposes. This output frequency is selected as it adequately captures relevant flow dynamics while significantly reducing the number of temporal snapshots. This approach enhances computational efficiency for the subsequent AI model training, which can effectively learn the continuous dynamics from this sparser data, thus reducing overall training overhead. Importantly, these simulations feature no external forcing after $t=0$, resulting in gradually dissipating flow fields where vorticity magnitude decreases over time. This dissipative behavior creates challenging temporal dynamics that are particularly difficult to predict accurately as the flow evolves further from its initial state. Throughout this study, we focus on predicting the vorticity field, a derived quantity that captures the local rotation in the fluid and presents a more challenging prediction target than primary variables like velocity components. Figure 5 displays the vorticity field at $t=0.5s$ for four distinct initial conditions. The marked diversity in the flow patterns, resulting from variations in $var_{1}$ and $var_{2}$, underscores the complex nature of the dataset.
@@
Figure 5: Vorticity fields at $t=0.5s$ for four distinct initial conditions, each defined by a different combination of $var_{1}$ and $var_{2}$. This diversity in the observed patterns highlights the complexity and challenges of predicting the flow evolution.
For the LF datasets, we generate 300 training samples for each of three different grid resolutions: 16×16, 32×32, and 64×64. While the generation of HF 128×128 data requires approximately 360 seconds per sample using AMD EPYC 7282 processor, the LF data require 90, 100, and 140 seconds for 16×16, 32×32, and 64×64 resolutions, respectively. In total, generating the HF dataset (100 training samples) requires 10 hours, whereas generating each LF dataset (300 samples) requires 7.5, 8.3, and 11.7 hours, respectively. As illustrated in Figure 6, the grid resolution significantly impacts the accuracy of flow field predictions. The coarsest resolution (16×16) shows severe limitations in capturing flow evolution: while it approximates the macroscopic flow patterns at $t=0.25s$, the flow field becomes notably distorted from $t=0.5s$ onward compared to higher-resolution simulations. This suggests that data from such coarse resolutions might be inadequate or even detrimental in a multi-fidelity framework. However, from 32×32 resolution upward, the simulations begin to capture the essential macroscopic flow properties visible in the HF (128×128) case. This observation suggests that LF data from these intermediate resolutions (32×32 and 64×64) could effectively support HF predictions in our multi-fidelity framework.
Figure 6: Effects of grid resolution on vorticity field evolution with fixed parameters $var_{1}=3$ and $var_{2}=1$ (contours represent vorticity values). It also shows how dissipative effects intensify over time in the adopted flow field dataset.
Note on dataset nesting in multi-fidelity modeling: Although the LF datasets used in this study are spatially nested within the HF datasets—i.e., their grid points are subsets of the higher-resolution grids (e.g., 64×64 is fully contained within 128×128)—the input function parameters ($var_{1}$ and $var_{2}$) are sampled independently across fidelities. Consequently, the datasets are not nested in terms of input function parameters. This is a critical distinction because multi-fidelity strategies often assume corresponding LF and HF outputs exist for identical input queries [29, 9, 10], which is not the case here due to independent parameter sampling across fidelities. This design choice aligns with our objective to develop an MF-DeepONet framework that operates robustly without requiring paired or nested datasets, thereby enhancing its practical applicability in real-world scenarios where consistent parameter sampling across fidelities is often infeasible. It should also be noted that our framework can accommodate datasets with non-nested spatio-temporal grid points, which is not explored in this study.
@@ 4 Improvement of DeepONet Architecture for Spatio-Temporal Flow Field Prediction
In this section, we focus on enhancing single-fidelity DeepONet performance using only high-fidelity data (128×128 resolution). We systematically investigate the effects of different merge processes, spatial point subsampling approaches, and positional encoding strategies to establish the most effective base architecture. This optimized DeepONet architecture will serve as the foundation for both LF and HF models in our multi-fidelity framework presented later.
@@ 4.1 Effects of Different Merge Process of Branch/Trunk Nets
We first investigate the effectiveness of different strategies for combining outputs from the branch and trunk networks elaborated in Section 2.1.1. To ensure a fair comparison, we carefully adjust the network architectures to maintain similar numbers of learnable parameters across different configurations (approximately 300K parameters). For the basic dot-product variants, Merge Type 0 with and without bias, we employ more complex branch and trunk networks to match the parameter count of models with Type 1 and 2 networks. Additionally, we implement the modified DeepONet architecture proposed by Wang et al. [48], which introduces encoder layers for enhanced nonlinear interactions but still relies on the fundamental dot-product mechanism between branch and trunk networks at its core. Across all DeepONet configurations, the branch network consistently takes two input parameters ($var_{1}$ and $var_{2}$) defining initial conditions, while the trunk network processes three spatial-temporal coordinates ($x$, $y$, and $t$). For detailed network architectures, please refer to Table 1. All models are trained for 800 epochs using the Adam optimizer with an initial learning rate of $10^{-3}$, GeLU activation functions, and minibatch size of 1 on an NVIDIA A6000 GPU. To account for training stochasticity and ensure reproducible results, all experiments in this paper are repeated at least three times with different random seeds, and we report the averaged values of these runs. To evaluate the predictive accuracy of each model, we report the test error as the root mean squared error (RMSE) computed across all spatio-temporal test points and test cases. Specifically, for a set of $N_{\text{test}}$ test cases, each with $N_{t}$ temporal snapshots and $N_{x}$ spatial points, the RMSE is defined as:
$\text{RMSE}=\sqrt{\frac{1}{N_{\text{test}}\cdot N_{t}\cdot N_{x}}\sum_{i=1}^{N_{\text{test}}}\sum_{j=1}^{N_{t}}\sum_{k=1}^{N_{x}}\left(\hat{\omega}_{ijk}-\omega_{ijk}\right)^{2}}$ (7)
where $\hat{\omega}_{ijk}$ and $\omega_{ijk}$ denote the predicted and ground truth vorticity values, respectively, for the $i$-th test case, at time $t_{j}$, and spatial location $\mathbf{x}_{k}$. All test errors reported in this paper, including those in Table 1, are expressed as these RMSE values: this definition is expected to ensure that the reported test error reflects a comprehensive average across all spatio-temporal predictions in the test set, providing a physically meaningful and consistent measure of overall model accuracy.
@@
Wang et al. [48]
Branch: 2-128
Trunk: 3-64-128
Merge: 256-128×5-64-32-16-1
302,336 17.918 32.099
The results in Table 1 demonstrate three key findings. First, incorporating merge networks significantly improves prediction accuracy compared to the two Type 0 approaches suggested by Lu et al. [26]. While the original dot-product architecture was successful for simpler operator learning tasks in other studies, our results show its limitations in complex flow field predictions: the test error with dot-product operations (both with and without bias) remains above 34, indicating substantial prediction inaccuracies. In contrast, the introduction of merge networks dramatically reduces the test error to 17-23, despite requiring longer training times (approximately 8.5 hours compared to 6.2 hours) due to the additional processing of branch and trunk outputs through the distinct merge network. This stark improvement in accuracy—reducing test error from 34.447 to 17.079, a 50.4% improvement—highlights that additional non-linear processing of combined features is crucial for capturing complex spatio-temporal flow dynamics. The merge network effectively acts as a learnable non-linear mapping that can adapt to intricate relationships between the latent representations from branch and trunk networks, a capability that simple dot-product operations alone cannot provide.
Second, the comparison with Wang et al. [48]’s approach provides additional evidence for the limitations of dot-product mechanisms in complex flow field predictions. Despite having a similar parameter count (302,336 parameters), Wang et al. [48]’s method achieves a test error of 32.099, which is substantially higher than our Type 1 (17.079) and Type 2 (23.060) approaches. Moreover, this approach requires significantly longer training time (17.918 hours) compared to our Type 1 (8.672 hours) and Type 2 (8.399 hours) methods, indicating computational inefficiency due to its additional architectural complexity. Notably, while this approach outperforms the basic dot-product operations (Type 0), it still falls short of our proposed merge network strategies in both accuracy and computational efficiency.
Third, among the enhanced architectures, Type 1 (element-wise multiplication) outperforms Type 2 (concatenation), achieving a test error of 17.079 compared to 23.060. This advantage stems from the structured interaction in Type 1, which more naturally aligns with well-established techniques such as modal decomposition (e.g., proper orthogonal decomposition) and spectral expansions, both of which have demonstrated strong performance in CFD.
@@
    Spectral Methods Perspective: Type 1 approach also aligns with spectral methods, where solutions are represented as weighted sums of basis functions [24]. Here, the trunk network acts as a global basis generator, while the branch network provides coefficients—akin to spectral expansions in Fourier or Laplace methods. Since spectral methods have already demonstrated strong performance in solving the Navier-Stokes equations [4, 40], enforcing this structured decomposition allows Type 1 to better capture the underlying physics of flow evolution, outperforming the less structured feature fusion in Type 2.
@@
Note on overfitting issues in over-parameterized DeepONet: To obtain the architectures summarized in Table 1, we conducted manual hyperparameter tuning across a wide range of model depths, hidden layer sizes, and merge configurations with varying parameter numbers. Despite testing deeper and wider networks with significantly larger parameter counts, we consistently observed that such models led to worse generalization performance—even as they achieved lower training losses. This suggests a critical insight: DeepONet tends to overfit easily when over-parameterized, indicating that its expressive power must be tightly controlled. The final architectures, each with approximately 300k parameters, were selected based on their superior balance between training efficiency and test accuracy.
4.2 Effects of Spatial Point Subsampling and Application of Automatic Mixed Precision
Building upon the optimal merge network configuration (Type 1) identified in Section 4.1, we investigate the impact of spatial point subsampling and automatic mixed precision (AMP) training on model performance. Spatial point subsampling refers to our approach of selecting only a subset of available spatial locations from the full 128×128 grid for training, leveraging DeepONet’s inherent capability to operate on arbitrary spatial points without requiring fixed grid structures. We explore three subsampling strategies: 16×16, 32×32, and 64×64 randomly selected points from the original 128×128 grid. The optimal choice of this subsampling ratio (e.g., the number of points like 16×16 or 32×32) can be highly problem-dependent, influenced by factors such as the physical characteristics of the investigated flow field, the scale of its primary dynamics, and overall data complexity. Therefore, for new applications, we recommend conducting a preliminary parametric study, similar to the exploration presented in Table 2, to determine a suitable subsampling level that effectively balances predictive accuracy with computational training efficiency.
AMP training refers to the strategic use of different numerical precisions (e.g., float32, float16) during model training to reduce memory usage and computational overhead without compromising model accuracy—a particularly relevant approach for DeepONet due to its intensive computational requirements in processing high-dimensional flow field data (for the details of AMP, please refer to Appendix A). For each subsampling configuration, we conduct experiments both with and without AMP training to systematically evaluate their combined effects on computational efficiency and prediction accuracy. All other training settings remain consistent with Section 4.1.
The results shown in Table 2 reveal important insights about DeepONet’s training efficiency through the combination of sampling strategies and AMP training. First, examining the impact of spatial point sampling, we observe that smaller sampling sizes can maintain prediction accuracy while significantly reducing computational costs. With AMP enabled, the 16×16 sampling strategy achieves comparable accuracy (test error: 17.419) to the baseline that uses all available points (test error: 17.079), while dramatically reducing the training time from 8.672 to 0.350 hours—a 96% reduction. Similar trends are observed in cases without AMP, where the 16×16 sampling rather achieves a test error of 16.876 compared to 18.001 for full sampling due to its effective sampling nature compared to naive all-point-sampling approach.
Table 2: Comparison of different spatial subsampling strategies with and without AMP training: Type 1 merge network is applied.